OpenStudy (anonymous):

There is 1.75 g of Al in an Aluminum Carbonate sample. What is the mass in g of Oxygen in the solution?

OpenStudy (anonymous):

\[Al _{2}(CO _{3})_{3}\]

OpenStudy (anonymous):

As you can see, each aluminum carbonate molecule contains 2 Al atoms and 9 O atoms. Therefore, to find the mass of O, you can use\[m _{O}=\frac{ n _{Al} }{ 2 }\times 9 \times M _{O}=\frac{ \frac{ 1.75 }{ 26.982 } }{ 2 }\times 9 \times 15.999= ...\]